# Scientific judgment drifts over time in AI ideation

### v1

@@ Abstract
Scientific discovery begins with ideas, yet evaluating early-stage research concepts is a subtle and subjective human judgment. As large language models (LLMs) are increasingly tasked with generating scientific hypotheses, most systems implicitly assume that scientists’ evaluations form a fixed gold standard, assuming that scientists’ judgments do not change. Here we challenge this assumption. In a two-wave study with 7,182 ratings from 57 active researchers across six scientific departments, each participant repeatedly evaluated a constant “control” research idea alongside AI-generated ideas. We show that scientists’ ratings of the very same idea systematically drift over time: overall quality scores increased by 0.61 points on a 0-10 scale (P = 0.005), and test–retest reliability was only moderate across core dimensions of scientific value, revealing systematic temporal drift in perceived idea quality. Yet the internal structure of judgment remained stable, such as the relative importance placed on originality, feasibility, clarity, and other criteria. We then aligned an LLM-based ideation system to first-wave human ratings and used it to select new ideas. Although alignment improved agreement with Wave-1 evaluations, its apparent gains disappeared once drift in human standards was accounted for. Thus, tuning to a fixed human snapshot produced improvements that were transient rather than persistent. These findings reveal that human evaluation of scientific ideas is not static but a dynamic process with stable priorities and requires shifting calibration. Treating one-time human ratings as immutable ground truth risks overstating progress in AI-assisted ideation and obscuring the challenge of co-evolving with changing expert standards. Drift-aware evaluation protocols and longitudinal benchmarks may therefore be essential for building AI systems that reliably augment, rather than overfit to, human scientific judgment.
Scientific ideation is the engine of technical progress (?, ?). Yet the ever-expanding body of literature (?, ?) and the rise of narrow specialization (?) have together created what Jones called the “burden of knowledge” (?): as mastery deepens, exploring new directions becomes harder. Automating the generation of research ideas offers a way to aggregate dispersed expertise and to probe a broader space of possibilities. Because of its intellectual complexity, automated ideation has long been regarded as a grand challenge for Artificial Intelligence (?, ?, ?).
@@
Evaluating the quality of generated ideas, however, remains a central obstacle. Implementing each idea and assessing them by outcomes is infeasible due to the resource- and time-intensive nature of scientific work. A practical alternative is to have scientists rate their potential value. However, designing evaluations to reflect genuine scientific value is difficult. Some studies use LLMs as proxies for human evaluators, either calibrated on human ratings (?) or justified by their correlation with human scores (?, ?), while others rely entirely on expert judgment to guide iteration (?, ?, ?). In all cases, human ratings are treated as a fixed “gold standard” for evaluation. Ideation systems are optimized to match that snapshot, assuming that scientists’ criteria are stable and the resulting improvements will persist over time.
@@
The risk echoes Goodhart’s law: a measure ceases to be good once it becomes a target (?). History offers cautionary examples of what happens when human variability is ignored. In the 1990s, U.S. health systems adopted patients’ self-reported pain level as a core quality metric. As patients’ expectations of pain relief and reporting behavior changed, opioid prescriptions escalated and caused widespread harm (?, ?). Similar dynamics arise in LLM training. Models must balance factual accuracy and user agreement. When optimization focuses narrowly on feedback collected at a single time point, it can overshoot and amplify the tendency rewarded most strongly at that time. Recent examples include “sycophancy” in GPT-4o (?, ?), where models learned to agree excessively with users after alignment updates based on static preference data, followed by subsequent over-corrections once the issue was recognized (?). In both medicine and AI, optimizing against an unmodeled, time-dependent human metric has produced fragile and unintended outcomes.
@@
We conducted a two-wave, within-participant study (Fig. 1). In each wave, every scientist rated six AI-generated ideas and one repeated human-written control idea (Fig. 1A). The repeated control idea provided a direct measure of temporal drift and test-retest reliability in ratings. For each idea, participants rated eight scientific value dimensions adapted from Dean et al. (?) and an Overall Quality score (0-10; 0=poor, 10=great; Fig. 1B). To model how scientists weighted dimensions when forming overall judgments, we regressed Overall Quality on the eight dimension scores. We tested whether these weights changed via Dimension $\times$ Wave interactions in a mixed-effects model. Between waves, we tuned the ideation system with Wave 1 human ratings. We then used the aligned evaluator to select Wave 2 AI ideas (Fig. 1C). We asked whether Wave 2 AI ideas were perceived as superior by the participants while controlling for drift using the unchanged control idea (Fig. 1D). This design enabled three analyses: (i) quantifying temporal changes in dimension-level perceptions and overall leniency, (ii) characterizing stability in cross-dimension weights, and (iii) testing whether tuning to a fixed human snapshot produces lasting gains while controlling for human judgment drift.
@@
Sixty active researchers from six departments at Duke University participated (21 faculty; 39 postdoctoral fellows, graduate students, or research staff; demographics and selection criteria in SM 1). Topics were customized to each participant’s domain to ensure subject familiarity (Methods 4.1). Participants were unaware that one idea would be repeated across waves. After excluding three outliers for extreme inconsistency (SM  8), the final sample comprised of $N=57$ participants, yielding a total of 7,182 individual ratings (57 participants $\times$ 2 waves $\times$ 7 ideas $\times$ 9 scales).
@@
Participants rated each idea on eight dimensions: Originality, Paradigm Relatedness, Acceptability, Implementability, Applicability, Effectiveness, Implicational Explicitness, and Clarity; all on ordinal scales (SM  13). Constructs and scales followed Dean et al. (?). Participants also rated on Overall Quality (0–10 scale).
To test whether perceptions of ideas changed over time, we implemented a test-retest design. Each participant re-evaluated the same human-written control idea in both waves. For each participant, the control idea was derived from a recent paper closely related to their field (excluding their own work; selection criteria in SM 5). We shuffled the appearing orders of the ideas in all surveys. Depending on the response time, inter-wave intervals ranged from days to weeks (median [IQR] = 24 [8-35] days; SM 9).
We tested mean change of the control idea in Overall Quality with a paired $t$-test and visualized agreement using a Bland–Altman plot. Dimension-level shifts were evaluated with Wilcoxon signed-rank tests and Benjamini–Hochberg false-discovery-rate (FDR) correction across eight comparisons. Test–retest reliability was estimated using intraclass correlation coefficient ICC(3,1) for Overall Quality and quadratic weighted kappa for each dimension.
@@
Expertise Factulty/Non-faculty
@@
We next tested whether the relative importance of the eight dimensions shifted. Verbal reports of evaluation criteria often diverge from the implicit rules guiding decisions (?). To infer the latent weighting of each dimension, we modeled the Overall Quality as a linear combination of the eight dimension scores across all ideas (one control and six AI ideas per participant) and both waves: Overall $\sim$ Dimensions + Wave + (Dimensions $\times$ Wave). All dimensions were $z$-scored within wave to enable coefficient comparability. We fit a mixed effects model with participant random intercepts and used the Wave $\times$ Dimension interaction terms to capture shifts in the relative importance of dimensions. BH-FDR correction was applied across the eight tests.
@@
Finally, we studied whether improvements achieved by tuning an ideation system to a fixed snapshot of human criteria persisted once evaluation standards changed. The system implemented a representative LLM-based ideation pipeline aligned to Wave 1 human ratings. Wave 2 AI ideas were then selected by this aligned evaluator and rated by the same participants. We tested whether Wave 2 AI idea ratings improved over Wave 1 after accounting for concurrent shifts in human standards, as quantified by the control idea’s change in rating. Next, we detail the core components of this study.
Capturing a snapshot of human criteria. LLM-driven idea generators are typically designed by grounding new ideas in prior literature (?, ?, ?, ?, ?, ?, ?). Following this principle, our system synthesized a pool of 100 candidate ideas per participant from structured summaries of their own publications (Methods 3). Since LLM-generated ideas vary in quality, previous systems often apply an automated evaluation step to score, refine or filter ideas (?). We again adopted this common practice by asking an LLM evaluator to rate each candidate idea using the same eight dimensions and prompts as the human raters. The Overall Quality score was computed as a weighted sum of dimensions. Without calibration, such LLM evaluators can diverge from human judgment both in per-dimension ratings and in inter-dimensional weighting. In Wave 1, we deployed the default evaluator with heuristic weighting to rank ideas and selected the top six for human evaluation (Methods 4.1). Prompts and LLM model specifics are provided in SM 12 and  13.
Aligning system evaluation criteria with human snapshot. After collecting the Wave 1 human ratings, we aligned the LLM evaluator to human scores in two steps: (i) fitting per-dimension linear calibrations mapping LLM scores to human scores and (ii) trained a ridge regression from the calibrated dimensions to the human-rated Overall Quality (train / validation split by idea; Methods 7). The fitted model represented a fixed snapshot of Wave 1 human criteria. For each participant, this aligned evaluator re-scored the same candidate pool and selected the six highest-rated ideas for Wave 2 evaluation. We validated the calibration in two ways. We compared the residual error on a validation set of ideas with a uniform weights model. We also compared the Pearson correlation of uncalibrated and calibrated LLM evaluator ratings with the human ratings.
@@ Results
Figure 2: Scientists’ evaluations of the same scientific idea drift over time. (A) Test-retest reliability of Overall Quality score for the repeated human-written control idea was moderate, indicating non-trivial variability even for identical stimuli. (B) Bland-Altman analysis of Overall Quality shows a systematic upward shift in ratings between waves with no evidence of proportional bias. (C) Test-retest reliability of 8 scientific value dimensions (quadratic weighted kappa with 95% CI). (D) Distribution of within-participant changes (Wave 2 – Wave 1) for the control idea shows a significant group-level increase (mean = +0.61, P = 0.005), while most individual changes fall within a wide variability range, consistent with subtle but systematic drift. (E) Paired trajectories for each participant across both waves illustrate consistent upward movement in Overall Quality and small positive shifts across most dimensions.
@@
Systematic drift: When the same participant re-rated the same control idea in Wave 2, the Overall Quality score increased by 0.61 points on a 0-10 scale (paired $t$ test, N = 57 participants, 95% CI $[0.19,1.04]$, P = 0.005, dz = 0.386; Fig. 2A). A Bland-Altman analysis showed positive bias of 0.61, with 95% limits of agreement from -2.50 to 3.73 and no evidence of proportional bias, i.e. higher-rated ideas did not drift differently from low-rated ones (slope = -0.125, P = 0.225, Fig. 2B).
@@
At the dimension level, all eight scientific value dimensions (six four-point scales and two three-point scales) shifted slightly upward, although none reached statistical significance in Wilcoxon signed-rank tests (all FDR-adjusted Ps $\geq$ 0.61; Fig. 2C). These results indicate a small but systematic upward drift in perceived idea quality, distributed broadly across dimensions rather than concentrated in any single criterion (Fig. 2E).
@@
Test-retest reliability: Overall Quality ratings showed moderate test-retest reliability (ICC(3,1) = 0.721, 95% CI [0.556, 0.826]). The Standard Error of Measurement (SEM) (1.124 points) yielded a Minimal Detectable Change of 3.116 points at 95% confidence (MDC95), suggesting that only shifts larger than $\sim$3.1 points can be confidently interpreted as true perceptual changes. Reliability across dimensions, measured by quadratic weighted kappa, was similarly moderate and heterogeneous: Originality 0.54 [0.33, 0.70], Implementability 0.57 [0.27, 0.79], Acceptability 0.64 [0.40, 0.79], Effectiveness 0.53 [0.30, 0.70], Applicability 0.49 [0.20, 0.71], Paradigm relatedness 0.51 [0.31, 0.65], with lower agreement for Implicational explicitness 0.37 [0.11, 0.57] and Clarity 0.40 [0.11, 0.61] (Fig. 2B). Although individual-level changes were typically smaller than the 3.1-point detection threshold, these shifts occurred in the same direction across most participants as shown in the last section, showing a statistically reliable group-level increase in perceived idea quality.
@@
Drivers of the shift: We modeled the control idea’s Overall Quality with participant random intercepts and fixed effects for Wave, Position, Context, Time gap, Time of day, Expertise, and S-wave (participants’ rubric-weighted perception), with pre-specified Wave $\times$ Context, Wave $\times$ Time gap , and Wave $\times$ S-wave interactions (Table 1). Results are shown in Fig. 3A. Ratings were higher at Wave 2 ($\beta_{\textrm{wave}}$ = 0.449, SE = 0.163, z = 2.752, P = 0.006), consistent with the paired-test results. S-wave strongly predicted Overall Quality ($\beta_{\textrm{S-wave}}$ = 1.727, SE = 0.165, z = 10.488, P << 0.001), indicating that participants’ stated evaluation rubrics tracked their actual judgments. Position, Context, Time gap, Time of day, and Expertise were not significant (Ps $\geq$ 0.30). Interactions with Wave were not significant (Ps $\geq$ 0.42; full results see Tab. S3) either. Thus, our results show that the observed drift cannot be explained by the experiment artifacts-related factors such as survey order or timing effects, suggesting that human criteria of Overall Quality themselves subtly evolve over time. Interestingly, the positive Wave effect, after controlling for S-wave, suggest that there are remaining variance in the score drift of the Overall Quality that cannot be explained by current predictors.
@@
We next asked whether scientists altered the relative importance assigned to each of the eight dimension. Across all ideas, we fit a mixed-effects model of Overall Quality as: Overall $\sim$ z-dimensions + Wave + (z-dimensions $\times$ Wave). Results are shown in Fig. 3B. Group-level coefficients showed strong positive associations for Effectiveness ($\beta$ = 0.583, SE = 0.088, z = 6.592, P << 0.001), Clarity ($\beta$ = 0.542, SE = 0.078, z = 6.935, P << 0.001), Originality ($\beta$ = 0.525, SE = 0.073, z = 7.205, P << 0.001), Implicational explicitness ($\beta$ = 0.492, SE = 0.083, z = 5.90, P << 0.001), Applicability ($\beta$ = 0.415, SE = 0.093, z = 4.462, P << 0.001) and Implementability ($\beta$ = 0.203, SE = 0.081, z = 2.506, P = 0.012), and non-significant association for Acceptability ($\beta$ = 0.105, P = 0.157) and Paradigm relatedness ($\beta$ = -0.031, P = 0.670).
Critically, we find no significant evidence that the weights applied to dimensions changed across waves (FDR-adjusted Ps $\geq$ 0.212 for all Dimension x Wave interaction terms; Fig. 3C). Hence, while participants became slightly more lenient overall, the cognitive weighting of dimensions of what makes a “good idea” remained stable over the study interval.
Figure 3: Determinants and structure of scientific judgment across time. (A) Mixed-effects regression examining factors associated with changes in ratings of the unchanged control idea. The only significant predictor was Wave, indicating a systematic upward drift in perceived idea quality not explained by order, contextual comparison, time gap, time of day, or expertise. Self-reported evaluation weights (S-wave) strongly predicted ratings, confirming internal consistency in participants’ stated rubrics. (B) Mixed-effects model of Overall Quality predicted by 8 dimensions, revealing the hierarchy of criteria driving scientific evaluation. (C) Wave $\times$ Dimension interaction terms reveal no significant changes in dimension weights over time, indicating that while absolute ratings drifted, the underlying structure of scientific judgment remained stable.
@@
Validating the alignment: Aligning the LLM evaluator to Wave 1 human ratings sharply reduced prediction error on a held-out validation set (1.86 vs. 15.35 of uniform weights’ residual; Fig. 4A) and increased correlation with human scores of the Wave 1 ideas (Pearson r: 0.193 unaligned to 0.320 aligned; Fig. 4B, left). Alignment thus improved the agreement between LLM evaluators and the human raters by capturing a fixed snapshot of human criteria at Wave 1.
Figure 4: Alignment to a snapshot of human judgment yields transient rather than persistent gains. (A) Aligning the LLM evaluator to Wave 1 human ratings substantially reduced prediction error on both training and held-out validation ideas, compared with a uniform-weights baseline. (B) Alignment increased agreement with human evaluations in Wave 1, but this improvement weakened in Wave 2, indicating reduced alignment once human standards drifted. (C) Raw human ratings initially suggested improvement: Wave 2 AI ideas scored 0.28 points higher than Wave 1 ideas, but note that the repeated human control idea increased by 0.61 points over the same period, revealing concurrent upward drift. (D) Difference-in-differences after drift correction: no persistent gain beyond the control’s drift.
Snapshot gains do not generalize under drift.. When the same human raters later evaluated Wave 2 ideas, the Wave 1 human-aligned LLM evaluator’s correlation with human ratings became significantly weaker for Wave 2 ideas (Fig. 4B, right). While raw human ratings of newly selected Wave 2 AI ideas were 0.28 points higher than those of Wave 1 AI ideas, this improvement did not persist after adjusting for concurrent drift in the unchanged control idea. In an idea-level mixed-effects DiD, the Wave × Idea coefficient was -0.333 (P = 0.338; Fig. 4C). A confirmatory participant-level DiD gave the same conclusion (mean change in [AI - control] from Wave 1 to Wave 2 = -0.333; 95% CI [-0.861, 0.194]; dz = -0.168; t = -1.266; P = 0.211; Fig. 4D). Thus, the apparent gains achieved by aligning to a fixed snapshot of human criteria did not persist once human judgment drift was accounted for, highlighting that human-in-the-loop tuning may result in transient rather than persistent improvements.

### v2

@@ Abstract
[…]
@@
Evaluating the quality of generated ideas remains a central obstacle. Implementing each idea and assessing them by outcomes is infeasible due to the resource- and time-intensive nature of scientific work. A practical alternative is to have scientists rate their potential value. However, designing evaluations to reflect genuine scientific value is difficult. Some studies use LLMs as proxies for human evaluators, either calibrated on human ratings (?) or justified by their correlation with human scores (?, ?), while others rely entirely on expert judgment to guide iteration (?, ?, ?). In all cases, human ratings are treated as a fixed “gold standard” for evaluation. Ideation systems are optimized to match that snapshot, assuming that scientists’ criteria are stable and the resulting improvements will persist over time.
@@
The risk echoes Goodhart’s law: a measure ceases to be good once it becomes a target (?). Such dynamics arise in LLM training. Models must balance factual accuracy and user agreement. When optimization focuses narrowly on feedback collected at a single time point, it can overshoot and amplify the tendency rewarded most strongly at that time. Recent examples include “sycophancy” in GPT-4o (?, ?), where models learned to agree excessively with users after alignment updates based on static preference data, followed by subsequent over-corrections once the issue was recognized (?). In AI, optimizing against an unmodeled, time-dependent human metric has produced fragile and unintended outcomes.
@@
We conducted a two-wave, within-participant study (Fig. 1).
In each wave, every scientist rated six AI-generated ideas and one repeated human-written control idea (Fig. 1A). The repeated control idea provided a direct measure of temporal drift and test-retest reliability in ratings. For each idea, participants rated eight scientific value dimensions adapted from Dean et al. (?) and an Overall Quality score (0-10; 0=poor, 10=great; Fig. 1B). To model how scientists weighted dimensions when forming overall judgments, we regressed Overall Quality on the eight dimension scores. We tested whether these weights changed via Dimension $\times$ Wave interactions in a mixed-effects model. Between waves, we tuned the ideation system with Wave 1 human ratings. We then used the aligned evaluator to select Wave 2 AI ideas (Fig. 1C).
We asked whether Wave 2 AI ideas were perceived as superior by the participants while controlling for drift using the unchanged control idea (Fig. 1D). This design enabled three analyses: (i) quantifying temporal changes in dimension-level perceptions and overall leniency, (ii) characterizing stability in cross-dimension weights, and (iii) testing whether tuning to a fixed human snapshot produces lasting gains while controlling for human judgment drift.
@@
63 active researchers from six departments at Duke University participated (21 faculty; 42 postdoctoral fellows, graduate students, or research staff; demographics and selection criteria in SM 1). Topics were customized to each participant’s domain to ensure subject familiarity (Methods 4.1). Participants were unaware that one idea would be repeated across waves. The final sample comprised of $N=63$ participants, yielding a total of 7,938 individual ratings (63 participants $\times$ 2 waves $\times$ 7 ideas $\times$ 9 scales). Sensitivity analyses across outlier removal methods are reported in SM 8.
@@
Participants rated each idea on eight dimensions: Originality, Paradigm Relatedness, Acceptability, Implementability, Applicability, Effectiveness, Implicational Explicitness, and Clarity; all on ordinal scales (SM 14). Constructs and scales followed Dean et al. (?). Participants also rated on Overall Quality (0–10 scale).
To test whether perceptions of ideas changed over time, we implemented a test-retest design. Each participant re-evaluated the same human-written control idea in both waves. For each participant, the control idea was derived from a recent paper closely related to their field (excluding their own work; selection criteria in SM 5). We shuffled the appearing orders of the ideas in all surveys. Depending on the response time, inter-wave intervals ranged from days to weeks (median [IQR] = 23 [9-35]; SM 10).
We tested mean change of the control idea in Overall Quality with a paired $t$-test and visualized agreement using a Bland–Altman plot. Dimension-level shifts were evaluated with Wilcoxon signed-rank tests and Benjamini–Hochberg false-discovery-rate (FDR) correction across eight comparisons. Test–retest reliability was estimated using intraclass correlation coefficient ICC(A,1) for Overall Quality and quadratic weighted kappa for each dimension.
@@
Expertise Faculty/Non-faculty
@@
We next tested whether the relative importance of the eight dimensions shifted. Verbal reports of evaluation criteria often diverge from the implicit rules guiding decisions (?). To infer the latent weighting of each dimension, we modeled the Overall Quality ($Q_{ij}$) evaluated by participant $i$ for idea $j$ as a linear combination of the eight dimension scores across all ideas (one control and six AI ideas per participant) and both waves:
$Q_{ij}=\beta_{0}+u_{i}+\beta_{W}\text{Wave}_{ij}+\sum_{k=1}^{8}\beta_{k}X_{ijk}+\sum_{k=1}^{8}\beta_{Wk}(\text{Wave}_{ij}\times X_{ijk})+\epsilon_{ij}$ (1)
where $\beta_{0}$ is the fixed intercept, $u_{i}$ the random intercept for participant $i$, $\text{Wave}_{ij}$ the wave indicator and $X_{ijk}$ the score of the $k$-th scientific value dimension. The coefficients $\beta_{k}$ capture the main effect (relative weight) of each dimension while the interaction terms $\beta_{Wk}$ test for temporal shifts in those weights across waves. Finally, $\epsilon_{ij}$ represents the residual error. To enable direct comparability of coefficients, all eight dimension scores were transformed into standard scores within their respective waves.
@@
Finally, we studied whether improvements achieved by tuning an ideation system to a fixed snapshot of human criteria persisted once evaluation standards changed. Wave 1 and Wave 2 each contained six AI-generated ideas per participant, selected from the same underlying candidate pool of 100 ideas using different evaluator criteria (heuristic weights in Wave 1; Wave-1-human-aligned weights in Wave 2). The two sets were therefore generally distinct but could overlap in principle. Across all participants, 2.78 out of 6 ideas appeared in both waves on average. Overlap was not unexpected: high-quality ideas might rank highly under both heuristic criteria and under aligned criteria. We treated the two sets as separate outputs of their respective selection processes and reported a robustness check restricted to non-overlapping ideas in Results. The system implemented a representative LLM-based ideation pipeline aligned to Wave 1 human ratings. Wave 2 AI ideas were then selected by this aligned evaluator and rated by the same participants. We tested whether Wave 2 AI idea ratings improved over Wave 1 after accounting for concurrent shifts in human standards, as quantified by the control idea’s change in rating. Next, we detail the core components of this study.
Capturing a snapshot of human criteria. LLM-driven idea generators are typically designed by grounding new ideas in prior literature (?, ?, ?, ?, ?, ?, ?). Following this principle, our system synthesized a pool of 100 candidate ideas per participant from structured summaries of their own publications (Methods 3). Since LLM-generated ideas vary in quality, previous systems often apply an automated evaluation step to score, refine or filter ideas (?). We again adopted this common practice by asking an LLM evaluator to rate each candidate idea using the same eight dimensions and prompts as the human raters. The Overall Quality score was computed as a weighted sum of dimensions. Without calibration, such LLM evaluators can diverge from human judgment both in per-dimension ratings and in inter-dimensional weighting. In Wave 1, we deployed the default evaluator with heuristic weighting to rank ideas and selected the top six for human evaluation (Methods 4.1). Prompts and LLM model specifics are provided in SM 13 and  14.
Aligning system evaluation criteria with human snapshot. After collecting the Wave 1 human ratings, we aligned the LLM evaluator to human scores in two steps: (i) fitting per-dimension linear calibrations mapping LLM scores to human scores and (ii) trained a ridge regression from the calibrated dimensions to the human-rated Overall Quality (train / validation split by idea; Methods 7). The fitted model represented a fixed snapshot of Wave 1 human criteria. For each participant, this aligned evaluator re-scored the same candidate pool and selected the six highest-rated ideas for Wave 2 evaluation. We validated the calibration in two ways. We compared the residual error on a validation set of ideas with a uniform weights model.
@@ Results
Moderate reliability with a positive mean trend: When the same participant re-rated the same control idea in Wave 2, the Overall Quality score increased by an average of 0.44 points on a 0-10 scale (paired $t$ test, N = 63 participants, 95% CI [0.12, 1.01][-0.12,\;1.01], P = 0.121, dz = 0.198; Fig. 2D). Sensitivity analysis confirms consistently positive mean shift (range +0.31+0.31 to +0.48+0.48 points), although its statistical significance depended on outlier handling method ( Table S1 in SM 8). A Bland-Altman analysis showed positive bias and no evidence of proportional bias, i.e. higher-rated ideas did not drift differently from low-rated ones (slope = -0.008, P = 0.862, Fig. 2B).
@@
At the dimension level, all scientific value dimensions except Applicability (six four-point scales and two three-point scales) shifted slightly upward, although none reached statistical significance in Wilcoxon signed-rank tests (all FDR-adjusted Ps $\geq$ 0.37; Table S3). These results indicate a small positive mean trend in perceived idea quality, distributed broadly across dimensions rather than concentrated in any single criterion (Fig. 2E).
Figure 2: Scientists’ evaluations of the same scientific idea drift over time. (A) Test-retest reliability of Overall Quality score for the repeated human-written control idea was moderate, indicating non-trivial variability even for identical stimuli. (B) Bland-Altman analysis of Overall Quality shows a positive mean shift between waves with no evidence of proportional bias. (C) Test-retest reliability of 8 scientific value dimensions (quadratic weighted kappa with 95% CI). (D) Distribution of within-participant changes (Wave 2 – Wave 1) for the control idea. Mean change was positive for Overall Quality and all dimensions except Applicability; statistical significance varied with outlier handling (see sensitivity analysis in SM 8). (E) Paired trajectories for each participant across both waves illustrate consistent upward movement in Overall Quality and small positive shifts across most dimensions.
@@
Test-retest reliability: Overall Quality ratings showed moderate test-retest reliability (ICC(A,1) = 0.586, 95% CI [0.348, 0.759]; Fig. 2A). The Standard Error of Measurement (SEM) (1.585 points) yielded a Minimal Detectable Change of 4.394 points at 95% confidence (MDC95), suggesting that only individual shifts larger than $\sim$4.4 points can be confidently interpreted as true perceptual changes. Note that even when individual measurements are noisy, aggregating within-person differences across many participants can yield a reliable estimate of the group mean trend. Reliability across dimensions, measured by quadratic weighted kappa, was similarly moderate and heterogeneous: Originality 0.60 [0.40, 0.77], Implementability 0.52 [0.24, 0.74], Acceptability 0.63 [0.38, 0.78], Effectiveness 0.54 [0.33, 0.71], Applicability 0.42 [0.17, 0.65], Paradigm relatedness 0.56 [0.41, 0.69], with lower agreement for Implicational explicitness 0.29 [0.01, 0.52] and Clarity 0.28 [0.02, 0.54] (Fig. 2C). These results indicate that instability is not limited to the Overall Quality score, but extends across the component dimensions used to evaluate scientific ideas.
@@
Drivers of the shift: We modeled the control idea’s Overall Quality with participant random intercepts and fixed effects for Wave, Position, Context, Time gap, Time of day, Expertise, and S-wave (participants’ rubric-weighted perception), with pre-specified Wave $\times$ Context, Wave $\times$ Time gap, and Wave $\times$ S-wave interactions (Table 1). Results are shown in Fig. 3A. Wave coefficient was positive but not significant ($\beta_{\textrm{wave}}$ = 0.144, SE = 0.204, z = 0.706, P = 0.480), consistent with the positive but non-significant trend from the paired $t$ test. S-wave strongly predicted Overall Quality ($\beta_{\textrm{S-wave}}$ = 1.839, SE = 0.209, z = 8.778, P << 0.001), indicating that participants’ stated evaluation rubrics tracked their actual judgments. Position, Context, Time gap, Time of day, and Expertise were not significant (Ps $\geq$ 0.159). Interactions with Wave were not significant (Ps $\geq$ 0.289; full results see Tab. S4).
These predictors therefore do not explain the observed variability in repeated ratings. Several mechanisms remain plausible. One possibility is procedural familiarity: by Wave 2, participants may better understand the eight dimensions, reducing cognitive load and subtly changing how they score ideas. Our exploratory ablation analysis provides limited support for this explanation (SM 9). A second possibility is cross-wave contrast or adaptation: as participants are exposed to different AI-generated ideas over time, their implicit baseline for judging the repeated human-written control idea may shift. Note that within-wave contextual contrast (the mean rating of preceding ideas) was not a significant predictor in our model, but cross-wave adaptation cannot be fully isolated with our design.
Finally, broader changes in scientific perspective arising from reflection, reading, or evolving field norms may also contribute. Our study was designed primarily to detect and quantify the consistency of human rating. We leave further exploration of drift mechanisms for future work. Regardless of mechanism, the practical implication of our study remains consistent: one-time human ratings are a moving target and should not be treated as immutable ground truth for iterative alignment.
@@
We next asked whether scientists altered the relative importance assigned to each of the eight dimensions. Across all ideas, we fit a mixed-effects model specified in Eq. (1). Results are shown in Fig. 3B. Group-level coefficients showed strong positive associations for Applicability ($\beta$ = 0.554, SE = 0.096, z = 5.758, P <<0.001), Clarity ($\beta$ = 0.545, SE = 0.077, z = 7.059, P << 0.001), Implicational explicitness ($\beta$ = 0.502, SE = 0.084, z = 5.953, P << 0.001), Effectiveness ($\beta$ = 0.452, SE = 0.092, z = 4.917, P << 0.001), Originality ($\beta$ = 0.438, SE = 0.073, z = 5.999, P << 0.001), and Implementability ($\beta$ = 0.314, SE = 0.079, z = 4.000, P << 0.001), and non-significant association for Acceptability ($\beta$ = 0.118, P = 0.116) and Paradigm relatedness ($\beta$ = 0.022, P = 0.763).
Critically, we find no significant evidence that the weights applied to dimensions changed across waves (FDR-adjusted Ps $\geq$ 0.071 for all Dimension x Wave interaction terms; Fig. 3C). Hence, while repeated evaluations of the same idea exhibit variability over time overall, the cognitive weighting of dimensions of what makes a “good idea” remained stable over the study interval.
Figure 3: Determinants and structure of scientific judgment across time. (A) Mixed-effects regression examining factors associated with changes in ratings of the unchanged control idea. Self-reported evaluation weights (S-wave) strongly predicted ratings, confirming that participants’ stated rubrics align with their actual judgments. Other factors, including position, contextual comparison, time gap, time of day, and expertise, are not significant. (B) Mixed-effects model of Overall Quality predicted by 8 dimensions, revealing the hierarchy of criteria driving scientific evaluation. (C) Wave $\times$ Dimension interaction terms reveal no significant changes in dimension weights over time, indicating that while absolute ratings drifted, the underlying structure of scientific judgment remained stable.
@@
Validating the alignment: Aligning the LLM evaluator to Wave 1 human ratings sharply reduced prediction error on a held-out validation set (1.86 vs. 15.35 of uniform weights’ residual; Fig. 4A), capturing a fixed snapshot of human criteria at Wave 1.
Figure 4: Alignment to a snapshot of human judgment yields transient rather than persistent gains. (A) Aligning the LLM evaluator to Wave 1 human ratings substantially reduced prediction error on both training and held-out validation ideas, compared with a uniform-weights baseline. (B) Raw human ratings initially suggested improvement: Wave 2 AI ideas scored 0.29 points higher than Wave 1 ideas. The repeated human control idea increased by 0.44 points over the same period, revealing concurrent upward drift. (C) Difference-in-differences after drift correction: no persistent gain beyond the control’s drift.
Snapshot gains do not generalize under drift. While raw human ratings of newly selected Wave 2 AI ideas were 0.29 points higher than those of Wave 1 AI ideas, this improvement did not persist after adjusting for concurrent drift in the unchanged control idea. For each participant, we subtracted their own control idea change (Wave 2 - Wave 1) from their AI idea change - a participant-level difference-in-differences that assumes each individual drifts in their own way. We find that the AI idea rating improvement in Wave 2 disappears after adjusting for concurrent drift (mean = 0.156-0.156; 95% CI [0.766, 0.454][-0.766,\;0.454]; $d_{z}=-0.064$; t=0.511t=-0.511; $P=0.611$; Fig. 4C). As a robustness check, restricting the analysis to non-overlapping AI ideas (those uniquely contributed by each selection process) yielded the same conclusion: a raw improvement of +0.57+0.57 points did not persist after drift correction (+0.12+0.12 points, $P=0.703$). Thus, the apparent gains achieved by aligning to a fixed snapshot of human criteria did not persist once human judgment drift was accounted for, highlighting that human-in-the-loop tuning may result in transient rather than persistent improvements.
Across both waves, the human-written control ideas (derived from vetted published papers) received substantially higher Overall Quality ratings than the AI-generated ideas. We note that the control idea was deliberately selected from recent, cited publications and therefore represents a high bar.